# A Survey of Reinforcement Learning from Human Feedback

### v1

@@ 3.2.9 Feature Traces
While many approaches either rely on hard-coded features or learn a model entirely end-to-end, it is also possible to actively elicit new features from human feedback. Feature traces were proposed by [36] as an approach to actively learn new relevant features. This type of feedback relies on a human operator to demonstrate a behavior in which a certain feature of interest, such as the distance to a sensitive object, monotonically increases or decreases. They make it possible to extend the set of features once the current set can no longer adequately explain the human feedback supplied through another type of feedback. The defining characteristics of feature traces are that they are of descriptive (intent) and explicitly (explicitness) given in an active manner (involvement) for a single (arity) segment (granularity) on an instance-level abstraction.
@@
Feature traces are strongly related to inter-temporal preferences (Section 3.2.3) since both types rely on changes in feature- or reward values in the course of a single trajectory. [36] propose to learn from feature traces by leveraging a Bradley-Terry model to learn the feature values, similar to other approaches that use such a model to learn reward values. Similar to importance feedback, feature traces rely on another type of feedback to actually make use of the learned features and is, therefore, a purely supplementary form of feedback. For instance, [36] use intervention feedback to train a reward model on the set of features derived using feature traces.
3.2.10 Similarity Queries
@@
Similarity queries are a feedback type aimed at learning a representation conforming to a notion of similarity and difference in the trajectory space. That aim is closely aligned with that of feature queries, though the actual queries are more similar to comparisons. The queries consist of triples of trajectories, with one anchor and two alternatives, for which the human has to decide which pair is more similar. Responses to similarity queries are given on observed behavior (involvement) with ternary arity, descriptive intent, and explicit feedback, while the granularity and abstraction may vary. This type of feedback was first introduced by [34], who used it to learn representations for reward learning. Similar to feature traces and importance, this is a supplementary type of feedback.
@@ 3.3 Initializations
Initializations are most commonly given by examples of successful task completions, either in the form of terminal or goal states [278], expert demonstrations [80, 2], labeled demonstrations [70], or ranked demonstrations [40]. Since this is not the main focus of our survey, we refer to the literature on inverse RL [7] for further details.
@@ 4 Label Collection
In this section, we explain how preference data can be collected for training a reward model. We first overview how queries can be generated for a given query type, then how the query type itself can be selected, and finally discuss issues arising in such human-computer interaction.
@@ 4.1 Active Learning
Following the terminology used in Bayesian active learning, we call the criterion to measure the quality of a query acquisition function, which is used to select which queries are presented to a human labeler. We first review the various acquisition functions used in RLHF for query selection and then discuss the extension to the choice of feedback type.
4.1.1 Acquisition Function
@@
As the RL agent is trained and its learned policy changes, the trajectories it generates will naturally evolve. Most work directly uses the trajectories obtained during RL training for preferences learning. Alternatively, the agent can also generate trajectories specifically to be used for querying (not necessarily for RL training), possibly with a learned transition model to control the sampling cost (e.g., [216]). This kind of active generation of queries can possibly lead to more informative ones.
In order to efficiently learn a suitable reward model, the agent must generate and select queries (Line 5 from Algorithm 1) so that it can quickly learn a good strategy using those queries. This selection is performed via a criterion, usually called acquisition function, which allows the queries to be compared. An efficient acquisition function may need to include various factors, such as: uncertainty, on-policy data, query simplicity, trajectory quality, query diversity, and query cost, which will be discussed one by one in the following. As a side remark, as noted by [104], interestingly, the queries asked by an RL agent also reveal its current reward learning stage.
@@
This corresponds to how uncertain the agent is about the ground-truth reward function. We are particularly interested in epistemic uncertainty in this setting, i.e., uncertainty that can be reduced by additional queries. This is usually one of the most important aspects to consider when deciding which query to ask. This uncertainty is usually either represented as a probability distribution (i.e., belief) in Bayesian approaches or using an ensemble of reward networks to approximate this belief.
With a belief representation, various classic Bayesian acquisition functions have been considered, such as the probability of improvement or the expected improvement (e.g., [64]). However, they require maintaining a distribution over reward functions (e.g., using Gaussian processes [64]) and, therefore, may not be suitable for more complex domains due to the computational complexity. Thus, simpler alternative criteria have been considered, such as information gain [185, 28] or variance. Although computationally heavy, volume removal (i.e., the minimum volume of the hypothesis set removed by an answer to a query) has also been studied [223, 19, 18].
@@
When using an ensemble instead of a direct belief representation, these criteria for epistemic uncertainty reduction correspond to measures of disagreement within the ensemble. Previous criteria could possibly be applied, but one popular candidate is the variance of the ensemble outputs [157, 180, 99].
@@
Recently, the average entropy of the ensemble outputs (assuming a Bradley-Terry model for the human answers) has also been used [157, 156, 203]. However, note that it does not quantify epistemic uncertainty but rather the uncertainty in the human’s answers, as provided by the response model of the human. Therefore, this criterion may not be suitable in the RLHF setting since it amounts to focusing on the queries for which an answer is expected to be the most random (according to the Bradley-Terry model). By definition of this model, the segments in the pairwise comparisons are the most similar in terms of returns and are, therefore, the hardest to answer for the human.
@@
In addition to epistemic uncertainty, one may also take utilities (i.e., returns or expected returns in RL) into account to select queries. In a Bayesian setting, this leads to acquisition functions such as expected value of information [189] or information gain over return differences [168], while in a non-Bayesian setting, the notion of regret, which measures the difference of performance between a policy optimal for the ground-truth reward function and a policy optimal for a learned reward function, can be used [268].
@@
Only focusing on uncertainty is likely insufficient or inefficient in RL. In particular, it may be important to favor more on-policy trajectories to guarantee the relevance of the generated queries. Indeed, improving reward learning in state-action regions that may never be visited with the current policy would lead to wasteful queries [168]. One simple approach to ensure that the data is more on-policy is by favoring more recently-generated trajectories [74].
@@
Selecting queries only based on their informativeness may lead to queries that are hard to answer for a human, which is, for example, the case for the average entropy. The ease of answering a query is important to alleviate the cognitive load of the human oracle. Some work specifically takes this aspect into account, for instance, by considering the similarity of consecutive queries [212] or the information gain. For this latter criterion, [30] show that in contrast to volume removal, it naturally leads to queries that are easier to answer for a human because information gain can be increased when the uncertainty in the human answer is lower.
@@
Most approaches directly use the trajectories generated during RL training. Especially early in training, these can be very bad with respect to the ground-truth reward function. In addition to that, they can be irrelevant or even contradictory for a given task [135]. Building queries on such trajectories may lead to unnatural queries for a human to respond to, such as comparing a very bad trajectory with an irrelevant one. [135] measure trajectory quality by optimizing over sampled reward functions.
@@
When asking many queries (in sequence or in batch), the diversity of the queries becomes especially crucial to avoid asking redundant queries. Most work follows a very myopic approach: Queries are often selected from a randomly generated set of potential queries, and sequences of queries are not really coordinated. Few work [31, 267] specifically tackles the former point, but the latter point is rarely considered, which may be due to its computational intractability. Indeed, planning ahead a sequence of queries would amount to solving a sequential decision-making problem under uncertainty over a combinatorial action space (i.e., the set of possible queries).
@@
The cost of generating queries may also be an important factor if the interaction of the human is live since it may not be practical to let the human wait before showing any queries [31]. In that case, it may be more important to quickly show some relatively good queries instead of computing the most informative ones.
Although this factor may not translate directly into an acquisition function, it may influence the choice of the acquisition function and its implementation in a given problem.
@@
Since various different acquisition functions have been considered, some effort [157, 156] has been made to compare them. Generally speaking, uncertainty-based criteria (e.g., variance or average entropy) seem to often perform better empirically compared to random selection, a query diversity-based criterion alone or combined with an uncertainty-based criterion. In addition, [74] empirically observed that a variance-based criterion performs better than a selection method only based on trajectory recency. Surprisingly, random selection has been shown to perform competitively in some cases [58, 119]. Thus, a better understanding of which acquisition function should be preferred in which situation or domain is still an open question.
@@
In addition, combinations of different criteria have naturally also been evaluated. For instance, [216] use four acquisition functions (high uncertainty, high novelty, high reward, low reward) in parallel. This approach has also been validated in a 3D environment [214]. A more sophisticated approach consists of considering a portfolio of acquisition functions and learning to select them using a multi-armed bandit approach [116].
@@
Various extensions to the basic setting have also been investigated. In the context of multiple human labelers, the issue of selecting reliable teachers to query arises [65]. Assuming all teachers have the same preferences, this can be modeled by incorporating a rationality coefficient $\beta$ into a Bradley-Terry model and estimating this factor:
$\max_{\theta}\prod_{i=1}^{N}\frac{1}{1+\exp(\beta({R}_{\psi}(\tau^{i}_{2})-{R}_{\psi}(\tau^{i}_{1})))}\,\text{,}$ (3)
@@
where a higher $\beta$ corresponds to a more reliable human (see Section 5.1.2). The setting in which this assumption does not hold, i.e., the labelers’ reward functions differ (a setting already considered in inverse RL [54]), has also been studied recently [281, 69, 188, 15]. Interestingly, a noisy oracle may provide more information than a completely reliable oracle because the frequency of erroneous answers given by the former is related to how much a segment is preferred to the other one [280, 48]. In contrast, only a binary preorder over segments can be inferred from the answers of a reliable and deterministic oracle, which may not be enough to recover the true reward function.

### v2

@@ 3.2.3 Representation-Specific Classes
While many approaches either rely on hard-coded features or learn a model entirely end-to-end, it is also possible to actively elicit new features from human feedback. Feature traces were proposed by Bobu et al. (2022) as an approach to actively learn new relevant features. This type of feedback relies on a human operator to demonstrate a behavior in which a certain feature of interest, such as the distance to a sensitive object, monotonically increases or decreases. They make it possible to extend the set of features once the current set can no longer adequately explain the human feedback supplied through another type of feedback. The defining characteristics of feature traces are that they are of descriptive (intent) and explicitly (explicitness) given in an active manner (involvement) for a single (arity) segment (granularity) on an instance-level abstraction.
@@
Feature traces are strongly related to inter-temporal preferences (Section 3.2.1) since both types rely on changes in feature or reward values in the course of a single trajectory. Bobu et al. (2022) propose to learn from feature traces by leveraging a Bradley-Terry model to learn the feature values, similar to other approaches that use such a model to learn reward values. Similar to importance feedback, feature traces rely on another type of feedback to actually make use of the learned features and is, therefore, a purely supplementary form of feedback. For instance, Bobu et al. (2022) use intervention feedback to train a reward model on the set of features derived using feature traces.
Similarity Queries
@@
Similarity queries are a feedback type aimed at learning a representation conforming to a notion of similarity and difference in the trajectory space. That aim is closely aligned with that of feature queries, though the actual queries are more similar to comparisons. The queries consist of triples of trajectories, with one anchor and two alternatives, for which the human has to decide which pair is more similar. Responses to similarity queries are given on observed behavior (involvement) with ternary arity, descriptive intent, and explicit feedback, while the granularity and abstraction may vary. This type of feedback was first introduced by Bobu et al. (2023), who used it to learn representations for reward learning.
@@ 3.3 Initializations
Initializations are most commonly given by examples of successful task completions, either in the form of terminal or goal states (Xie et al., 2018), expert demonstrations (Ibarz et al., 2018; Fu et al., 2018a; Palan et al., 2019; Lee et al., 2021b; Bıyık et al., 2022a; Abramson et al., 2022; Huang et al., 2023), labeled demonstrations (Du et al., 2023), or ranked demonstrations (Brown et al., 2019). It is even possible to infer some human preferences by the state of the environment alone, e.g., assuming the parts of the initial environment that are under the user’s control are largely set up in accordance with the user’s preferences (Shah et al., 2019; Lindner et al., 2021a). Since offline initialization is not the main focus of our survey, we do not cover these works in detail. We refer the interested reader to literature on inverse RL (Arora & Doshi, 2021) for further details on learning from demonstrations in particular.
3.4 Choice of Feedback Type
The best feedback type is not always clear and may depend on the task, the user, or the agent. It may also change over time as the agent learns more about the task. Table 4 may serve as a starting point to select a set of feedback types which may be applicable based on the possible user interaction and expertise, e.g., by the desired granularity, level of abstraction or involvement. This choice may be informed, among other features, by the task complexity or time horizon: Jain et al. (2015) compare purely passive (re-ranking based) with active (slight trajectory improvements) feedback and conclude that the former is usually preferred by humans when the action space is large while the latter is preferred for complex tasks. Similarly, Sumers et al. (2023) empirically demonstrate that language is a more effective teaching modality than demonstrations for complex concepts. The relevance of the time horizon is highlighted in the work by Sumers et al. (2022b), which suggests that instructive feedback is more efficient for short time horizons and descriptive feedback for longer horizons.
In addition to these static choices, Section 4.1.2 will discuss how to choose a feedback type adaptively. It is also possible to let the user choose the feedback type, as demonstrated by Jain et al. (2015) who let the user choose between re-ranking and improvement. Jeon et al. (2020) propose to use this choice of feedback type itself as a source of information (“meta-choice”).
3.5 Combination of Feedback Types
In addition to using any of the previously described feedback types in isolation, combining multiple feedback types is both possible and often advantageous. There are three main ways to combine feedback types: (a) a two-phase setup, consisting of initialization and refinement, (b) integrating a primary feedback type with a supplementary one; and (c) merging multiple primary feedback types.
The two-phase setup can be used to either initialize the reward model from offline data or to learn a representation that improves later reward learning. A common approach involves using one feedback type, typically demonstrations (see Section 3.3), to initialize the reward model, subsequently fine-tuning this model with another type, such as comparisons. This method is exemplified by Ibarz et al. (2018), who combined demonstrations and pairwise comparisons. For a more detailed discussion on feedback types suitable for initialization, refer to Section 3.3. Alternatively, a representation-focused type of feedback might be employed initially (Section 3.2.3) to cultivate a superior representation, followed by the application of a primary feedback type for reward model learning.
Combining a primary feedback type with a supplementary one can be beneficial to make the most out of the available user-interactions. Supplementary feedback, while not sufficient for reward model training by itself, can often be collected cheaply or as a side-effect of other interactions, making it a valuable addition to the primary feedback type. Refer to Section 3.2.2 for a discussion on supplementary feedback types.
Finally, combining multiple primary feedback types can be beneficial to capitalize on the strengths of each. For instance, Koppol et al. (2020) combine informative and demanding queries with less-informative and less-demanding ones to achieve a balance between cognitive load and informativeness. Similarly, to enhance expressivity, Mehta & Losey (2023) combines demonstrations, pairwise comparisons, and corrections, allowing users to select their preferred type of feedback.
@@ 4 Label Collection
[…]
Figure 3: Highlighting the components of the RLHF loop discussed in this section.
In this section, we explain how preference data can be collected for training a reward model, independent of the specific type of feedback used. We start by overviewing the active learning problem posed by query selection, e.g., how queries can be generated for a given query type, and how even the type of feedback itself can be selected. Following this, we discuss issues arising in such human-computer interaction.
@@ 4.1 Active Learning
We first discuss query generation and selection for a given feedback type, then the extension of these techniques to the choice of feedback types.
4.1.1 Query Generation and Selection
@@
As the RL agent is trained and its learned policy changes, the trajectories it generates will naturally evolve. Most work directly uses the trajectories obtained during RL training for preferences learning. Using such trajectories, candidate queries are often randomly generated. However, for pairwise comparison, a more efficient approach is to generate queries by exploiting preference transitivity (Hwang et al., 2023). Alternatively, the agent can also generate trajectories specifically to be used for querying (not necessarily for RL training), possibly with a learned transition model to limit the sampling cost (e.g., Reddy et al. (2020); Liu et al. (2023b)). This kind of active generation of queries can possibly lead to more informative ones.
Table 5: An overview of query selection strategies used in RLHF approaches.
References Factors
Uncertainty On-policy Query Trajectory Query Query
Data Simplicity Quality Diversity Cost
Daniel et al. (2014) Probability of improvement
Expected improvement
Upper confidence bound
Christiano et al. (2017) Ensemble variance
Sadigh et al. (2017) Volume removal
Bıyık et al. (2024) Volume removal
Wilde et al. (2018) Feasible space reduction
Ibarz et al. (2018) Random
Mindermann et al. (2018) Information gain
Cui & Niekum (2018) Information gain
Racca et al. (2019) Entropy
Bıyık et al. (2019) Mutual information
Bıyık et al. (2020) Information gain
Reddy et al. (2020) Ensemble averaged KL-
divergence to mean output
Novoseller et al. (2020) Dueling posterior sampling
Wilde et al. (2020; 2021) Maximum regret
Lee et al. (2021b) Ensemble-averaged entropy
Lindner et al. (2021b) Information gain
Katz et al. (2021) Posterior sampling
Myers et al. (2023) Expected value of information
Bıyık et al. (2024) Mutual information
Dwaracherla et al. (2024) Double Thompson sampling
Hu et al. (2024) Random
In order to efficiently learn a suitable reward model, the agent must generate and select queries (Line 5 from Algorithm 1) so that it can quickly learn a good strategy using those queries. This selection is performed via a criterion, usually called acquisition function [^1] We follow the terminology used in Bayesian active learning., which allows the queries to be compared. Although most existing work in RLHF uses an acquisition function that provides some measure of uncertainty (e.g., about the learned rewards), which is arguably one of the most important factors in active learning, an efficient acquisition function (see Table 5 for works dedicated to improving query selection) may need to include various additional other aspects, such as: on-policy data, query simplicity, trajectory quality, query diversity, or query cost, which will be discussed one by one in the following. As a side note, interestingly, as highlighted by Habibian et al. (2022), the queries asked by an RL agent may also reveal its current reward learning stage.
@@
This factor usually corresponds to epistemic uncertainty (Hüllermeier & Waegeman, 2021), which represents how uncertain the agent is about the ground-truth reward function. Epistemic uncertainty can be contrasted to aleatoric uncertainty, which describes inherent stochasticity in the system. While the latter cannot be fully eliminated, the former can be reduced with additional queries. Uncertainty is usually one of the most important aspects to consider when deciding which query to ask. It is usually represented either as a probability distribution (i.e., belief) in Bayesian approaches or using an ensemble of reward networks to approximate this belief. However, other representations are also possible, e.g., using the recently-proposed epistemic neural network (Osband et al., 2023).
With a belief representation, various classic acquisition functions have been considered in the Bayesian framework, such as the probability of improvement, the expected improvement, or the upper confidence bound. For instance, Daniel et al. (2014) compare those three criteria in a robotic domain and observe that the latter yield the best performance, but at a high cost in terms of number of queries, while the first asks the fewest number of queries, but with a slightly lower asymptotic performance. Other alternative criteria have been considered, such as volume removal (Sadigh et al., 2017; Basu et al., 2018; 2019) or information gain (Mindermann et al., 2018; Bıyık et al., 2019; 2020). The volume removal criterion uses the minimum volume of the hypothesis set removed by an answer to a query as the acquisition function and has been shown to be effective in practice. However, Bıyık et al. (2019) show that volume removal has uninformative global optima and argue that the practical effectiveness is due to non-convexity leading to local optima that are informative. They also show that optimizing for information gain has the same computational complexity while avoiding this flaw. One drawback of these Bayesian approaches is that they require maintaining a distribution over reward functions (e.g., using Gaussian processes (Daniel et al., 2014) or simpler probability distributions, such as Gaussian distribution, but with a linear reward model) and, therefore, may not be suitable for more complex domains due to the computational complexity or the strong assumption about the reward model.
@@
When using an ensemble instead of a direct belief representation, these criteria for epistemic uncertainty reduction correspond to measures of disagreement within the ensemble. Previous criteria could possibly be applied, but one popular candidate is the variance of the ensemble outputs (Lee et al., 2021b; Metcalf et al., 2023; Gleave & Irving, 2022) or equivalently its standard deviation (Eberhard et al., 2022).
@@
The average entropy of the ensemble outputs (e.g., when assuming a Bradley-Terry model for the human answers) has also been used (Lee et al., 2021b; a; Park et al., 2022). However, note that it does not quantify epistemic uncertainty but rather the aleatoric uncertainty in the human’s answers, as provided by the response model of the human. Therefore, this criterion may not be suitable in the RLHF setting since it amounts to focusing on the queries for which an answer is expected to be the most random (according to the Bradley-Terry model). By definition of this model, the segments in the pairwise comparisons are the most similar in terms of returns and are, therefore, the hardest to answer for the human.
In contrast to those acquisition functions that lead to deterministic query selection, sampling-based approaches have also been studied, from pure random selection (Ibarz et al., 2018) to Thompson sampling (Katz et al., 2021) and its variants (Novoseller et al., 2020; Dwaracherla et al., 2024). Recently, in the context of fine-tuning LLMs with pairwise comparison queries, Dwaracherla et al. (2024) shows that using the relatively novel epistemic neural network (Osband et al., 2023), double Thompson sampling (Wu & Liu, 2016), which naturally favors better elements to be compared, performs well experimentally.
@@
In addition to epistemic uncertainty, one may also take the outcomes into account to select queries, that is consider utilities (i.e., returns or expected returns in RL). In a Bayesian setting, this leads to acquisition functions such as expected value of information (Myers et al., 2023) or information gain over return differences (Lindner et al., 2021b), while in a non-Bayesian setting, the notion of regret, which measures the difference of performance between a policy optimal for the ground-truth reward function and a policy optimal for a learned reward function, can be used (Wilde et al., 2020).
Finally, it should be mentioned that naturally approaches with theoretical guarantees usually also use uncertainty as the main criterion for label collection (e.g., Novoseller et al. (2023), see Section 7 for more details).
@@
Only focusing on uncertainty is likely insufficient or inefficient in RL because the previous methods may focus on choosing queries to identify the reward function as precisely as possible uniformly on the whole state-action space. However, it may be important to favor more on-policy trajectories to guarantee the relevance of the generated queries for the current policy. Indeed, improving reward learning in state-action regions that may never be visited with the current policy would lead to wasteful queries (Lindner et al., 2021b). One simple approach to ensure that the data is more on-policy is by simply sampling from the current policy (Cui & Niekum, 2018) or favoring more recently-generated trajectories (Eberhard et al., 2022; Hu et al., 2024).
@@
Selecting queries only based on their informativeness may lead to queries that are hard for a human to answer, which is, for example, the case for the average entropy. The ease of answering a query is important to alleviate the cognitive load of the human oracle. Some work specifically takes this aspect into account, for instance, by considering the similarity of consecutive queries (Racca et al., 2019) or the information gain. For this latter criterion, Bıyık et al. (2019) show that in contrast to volume removal, it naturally leads to queries that are easier to answer for a human because information gain can be increased when the uncertainty in the human answer is lower.
@@
Most approaches directly use the trajectories generated during RL training. Especially early in training, these can be very bad with respect to the ground-truth reward function. In addition to that, they can be irrelevant or even contradictory for a given task (Katz et al., 2021). Building queries on such trajectories may lead to unnatural queries for a human to respond to, such as comparing a very bad trajectory with an irrelevant one. Katz et al. (2021) measure trajectory quality by optimizing over sampled reward functions. Similarly, Cui & Niekum (2018) generate trajectories using optimal policies for reward functions sampled from the current Bayesian belief.
@@
When asking many queries (in batch, in sequence), the diversity of the queries becomes especially crucial to avoid asking redundant queries. Most work (Christiano et al., 2017; Lee et al., 2021b; Verma & Metcalf, 2024) follows a very myopic approach: Queries are often selected from a usually randomly-generated set of potential queries, and sequences of queries are not really coordinated. While some work exists that specifically tackles the selection of a batch of diverse queries (Bıyık & Sadigh, 2018; Bıyık et al., 2024), the latter is rarely considered due to its computational intractability. Indeed, planning ahead a sequence of queries would amount to solving a sequential decision-making problem under uncertainty over a combinatorial action space (i.e., the set of possible queries). For diverse batch querying, previous work considered using clustering methods such as k-medoids (Bıyık & Sadigh, 2018) or more recently determinantal point processes, which define probability distributions that promote diversity (Bıyık et al., 2024).
@@
The cost of generating queries may also be an important factor if the interaction of the human is live since it may not be practical to let the human wait before showing any queries (Bıyık et al., 2024). In that case, it may be more important to quickly show some relatively good queries instead of computing the most informative ones. Although this factor may not translate directly into an acquisition function, it may influence the choice of the acquisition function and its implementation in a given problem.
@@
Since various different acquisition functions have been considered, some effort (Lee et al., 2021b; a) has been made to compare them. Generally speaking, uncertainty-based criteria (e.g., variance or average entropy) seem to often perform better empirically compared to random selection, a query diversity-based criterion alone or combined with an uncertainty-based criterion. In addition, Eberhard et al. (2022) empirically observe that a variance-based criterion performs better than a selection method only based on trajectory recency. Surprisingly, random selection has been shown to perform competitively in some cases (Christiano et al., 2017; Ibarz et al., 2018). Thus, a better understanding of which acquisition function should be preferred in which situation or domain is still an open question.
@@
In addition, combinations of different criteria have naturally also been evaluated. For instance, Reddy et al. (2020) use four acquisition functions (high uncertainty, high novelty, high reward, low reward) in parallel. This approach has also been validated in a 3D environment (Rahtz et al., 2022). A more sophisticated approach consists of considering a portfolio of acquisition functions and learning to select them using a multi-armed bandit approach (Hoffman et al., 2011).
@@
Various extensions to the basic setting have also been investigated. In the context of multiple human labelers, the issue of selecting reliable teachers to query arises (Daniels-Koch & Freedman, 2022). Assuming all teachers have the same preferences, this can be modeled for pairwise comparisons by incorporating a rationality coefficient $\beta$ into a Bradley-Terry model and estimating this factor:
$\max_{\psi}\prod_{i=1}^{N}\frac{1}{1+\exp(\beta({R}_{\psi}(\tau^{i}_{2})-{R}_{\psi}(\tau^{i}_{1})))}\,\text{,}$ (3)
@@
where a higher $\beta$ corresponds to a more reliable human (see Section 5.1.2). The setting in which this assumption does not hold, i.e., the labeler’s reward functions differ (a setting already considered in inverse RL (Choi & Kim, 2012)), has also been studied recently (Siththaranjan et al., 2024; Xue et al., 2023a; Dong et al., 2024; Myers et al., 2021; Bakker et al., 2022). Interestingly, a noisy oracle may sometimes provide more information than a completely reliable oracle. For instance, in Eq. 3, the frequency of erroneous answers given by the noisy oracle is related to how much a segment is preferred to the other one (Xu et al., 2020; Chan et al., 2021). In contrast, only a binary preorder over segments can be inferred from the answers of a reliable and deterministic oracle, which may not be enough to recover the true reward function.
Another notable recent work is that by Ellis et al. (2024) who raise the issue of identifiability of the ground-truth reward function: many reward functions result in the same optimal behaviors. The authors propose a framework that enables the generation of acquisition function for various definitions of reward similarity, such as the one discussed in Section 5.3.